### v2

# jina-vlm: Small Multilingual Vision Language Model

Abstract
We present jina-vlm, a 2.4B parameter vision-language model that achieves state-of-the-art multilingual visual question answering among open 2B-scale VLMs. The model couples a SigLIP2 vision encoder with a Qwen3 language backbone through an attention-pooling connector that enables token-efficient processing of arbitrary-resolution images. The model achieves leading results on standard VQA benchmarks and multilingual evaluations while preserving competitive text-only performance. Model weights and code are publicly released at https://huggingface.co/jinaai/jina-vlm.
1 Introduction
Vision-language models (VLMs) combine pretrained vision encoders with large language models to tackle tasks requiring joint visual and textual understanding (flamingo; llava). Recent VLMs have achieved strong results on visual question answering (VQA), OCR, and multimodal reasoning. However, two challenges limit their practical deployment. First, multilingual capabilities often degrade during vision adaptation: models that perform well on English benchmarks show uneven results across other languages (vlmsurvey). Second, high-quality VLMs remain computationally expensive to train and deploy, limiting accessibility for researchers and practitioners with constrained resources.
This work introduces jina-vlm, a 2.4B parameter VLM that addresses both challenges. The model aligns a SigLIP2-So400M/14-384 vision encoder (siglip2) with Qwen3-1.7B-Base (qwen3) through an attention-pooling connector, trained with a two-stage pipeline that explicitly incorporates multilingual data. Among open 2B-scale VLMs, jina-vlm achieves state-of-the-art performance on multilingual multimodal benchmarks including MMMB and Multilingual MMBench, demonstrating that small models can excel at cross-lingual visual understanding without sacrificing general capabilities. On standard English benchmarks spanning diagrams, charts, documents, and OCR, jina-vlm achieves the highest average score (72.3) across eight VQA benchmarks among 2B-scale VLMs. These results are enabled by two technical contributions: an efficient arbitrary-resolution pipeline that combines overlapping tiling with attention-based token pooling to reduce visual token count by 4, and a training recipe that incorporates text-only data to preserve the language understanding performance of the backbone LLM.
2 Related Work
VLM architecture and training. Modern VLMs follow an architecture introduced by PaLI (pali): a pretrained vision encoder extracts visual features, a connector projects them into the language model’s embedding space, and a decoder-only language model generates text conditioned on these visual tokens. Vision Transformers (ViTs) (vit) produce patch-level representations that the language model processes alongside text embeddings. This design is adopted by LLaVA (llava; llava_1_5; llava_uhd; llava_next_interleave; llava_onevision), QwenVL (qwen_vl; qwen_2_vl; qwen_2_5_vl), InternVL (internvl; internvl_1_5; internvl_2_5; internvl_3; internvl_3_5), and Ovis (ovis; ovis_2_5). Training strategies vary: qwen_2_vl; internvl_2_5 alternate between multimodal instruction tuning and general training; llava_1_5 incorporate academic VQA datasets; molmo, llava_onevision, and cambrian1 curate large-scale, diverse data mixtures.
Efficient resolution-agnostic image processing. Standard ViTs process fixed-resolution images, requiring resizing that discards fine-grained detail. Since visual token count scales with resolution and Transformer computation scales quadratically with sequence length, naive high-resolution processing is prohibitive. Several solutions exist: molmo tile images with overlap; qwen_2_vl introduce Naive Dynamic Resolution with Multimodal Rotary Position Embedding (rope; 2drope); ovis_2_5 use native-resolution ViTs (navit). Orthogonally, images often contain low-information regions (e.g., sky backgrounds), making visual tokens highly redundant. Token compression methods address this (fastv; prumerge; visionzip; pyramiddrop). internvl_1_5 develop Dynamic High-Resolution Tiling, and nvila propose scale-then-compress strategies. Recent work on training-free token budgeting, such as HERO (hero), demonstrates that inference-time pruning can achieve significant speedups while preserving accuracy; our approach differs by learning compact representations during training rather than dropping tokens at inference.
Vision-language connectors. The connector bridging vision encoders and language models significantly impacts both efficiency and performance. BLIP-2 (blip2) introduces Q-Former, a learnable query-based transformer that extracts fixed-length representations from visual features, reducing the number of tokens fed to the LLM. Flamingo (flamingo) uses a Perceiver Resampler with cross-attention to compress visual tokens. Our attention-pooling connector shares the goal of token reduction but operates differently: rather than learning a fixed set of queries, we apply local 22 attention pooling that preserves spatial structure while achieving 4 compression, which we found more effective for tasks requiring fine-grained spatial understanding.
Small VLMs. Efficiency has become a central objective. mobilevlmv2 demonstrate competitive performance below 2B parameters. imp combine quantization with aggressive resolution reduction for mobile deployment, matching larger models’ performance. MiniCPM-V (minicpmv) targets edge deployment while maintaining strong OCR and multilingual capabilities. smolvlm systematically explore design parameters to train VLMs as small as 256M parameters.
Multilingual VLMs. Many lightweight VLMs (paligemma; paligemma2; phi3) achieve strong English performance but degrade on other languages. qwen_2_vl and internvl_1_5 address this through targeted multilingual training data. pangea introduce instruction-tuning data spanning 39 languages.
Retaining text-only performance. Multimodal training often degrades text-only capabilities. Mitigation strategies include balanced data mixtures, careful learning rate scheduling (cauldron), and partial backbone freezing (llava_onevision; internvl_3_5).
3 Model Architecture
Figure 1 illustrates the architecture of jina-vlm. The model uses overlapping image tiling following molmo, combined with attention-based token pooling to reduce sequence length while preserving spatial information.
The vision encoder, SigLIP2-So400M/14-384, is a 27-layer Vision Transformer with 400M parameters that processes 378378 pixel inputs as 2727 grids of 1414 patches. To handle arbitrary resolutions, we decompose each image into overlapping tiles of this size and process each tile independently through the encoder. A global thumbnail, the full image resized to 378378, provides context alongside the tile representations. We use a default of 12 tiles during training; this limit can be increased at inference or during continued training to handle higher resolutions, with memory scaling linearly with tile count. The tiling algorithm is detailed in Appendix A.1.
3.1 Vision-Language Connector
Rather than using the final ViT output, jina-vlm concatenates features from two intermediate layers: the third-to-last and ninth-to-last, corresponding to layers 24 and 18 of the 27-layer encoder. This captures both fine-grained spatial details from earlier layers and high-level semantics from later layers. The connector then applies attention pooling over 22 patch neighborhoods, using mean-pooled features as queries. This reduces the token count by 4 while preserving local structure. A SwiGLU projection layer maps the pooled representations to the language model’s embedding dimension.
In more formal terms, let denote the hidden states from ViT layer , where is the number of patches, is the vision encoder hidden size, and negative indices count from the final layer (e.g., is the last layer). We concatenate features from two internal layers:
For each patch neighborhood , we compute a query vector as the mean of the neighborhood features:
where contains the four patches at positions , , , and and .
Attention pooling is then computed as:
where and , , and are learnable weight matrices. Finally, the pooled visual features are projected to the language model embedding dimension via a SwiGLU (swiglu) layer:
where , is the sigmoid function, denotes element-wise multiplication, , are learnable parameters, and is the language model embedding size.
3.2 Language Decoder
The language decoder is initialized from Qwen3-1.7B-Base111https://huggingface.co/Qwen/Qwen3-1.7B-Base, which empirically outperformed the instruction-tuned variant in our setting. We introduce three special tokens to structure visual inputs: <im_start> and <im_end> delimit image and thumbnail sequences, while <im_col> marks row boundaries within the patch grid, where tokens are arranged left-to-right and top-to-bottom. Input and output embedding weights are not tied.
3.3 Efficiency Analysis
Table 1 quantifies the computational benefits of attention pooling. With the default 12-tile configuration (plus thumbnail), the unpooled baseline would produce 9,477 visual tokens per image, while our 22 pooling reduces this to 2,366 tokens. Since the ViT processes each tile identically regardless of pooling, the savings apply exclusively to the LLM: we observe a 3.9 reduction in prefill FLOPs and a 4 reduction in KV-cache memory. The overall FLOPs reduction is 2.3 when including the shared ViT cost.
| Metric | No Pooling | With Pooling | Reduction |
| Visual tokens | 9,477 | 2,366 | 4.0 |
| LLM prefill FLOPs | 27.2 TFLOPs | 6.9 TFLOPs | 3.9 |
| KV-cache memory | 2.12 GB | 0.53 GB | 4.0 |
4 Training
Training proceeds in two stages, both updating all model components (encoder, connector, and decoder) without freezing, following molmo. The combined data comprises approximately 5M multimodal samples and 12B text tokens across 30+ languages, with roughly half in English and the remainder spanning high- and moderate-resource languages. Table 2 summarizes hyperparameters for both stages.
4.1 Stage 1: Alignment Training
The first stage focuses on cross-language semantic grounding rather than task-specific objectives. Training data consists primarily of caption datasets (PixmoCap (molmo), PangeaIns (pangea)) spanning diverse visual domains: natural scenes, documents, infographics, and diagrams. We include 15% text-only data from PleiAS/common_corpus (pleias) to mitigate degradation on text-only tasks. The connector uses a higher learning rate and shorter warmup than the encoder and decoder.
4.2 Stage 2: Instruction Fine-tuning
The second stage trains instruction-following for VQA and reasoning tasks. We combine public dataset collections, including LLaVA OneVision (llava_onevision), Cauldron (cauldron), Cambrian (cambrian1), PangeaIns (pangea), and FineVision (finevision), with text-only instruction data from aya. The mixture covers academic VQA, document understanding, OCR, mathematics, and reasoning. Appendix A.2 shows representative examples.
Given the diversity of instruction data, we found single-source batches more effective initially, likely due to the heterogeneous data mixture. We train for 30K steps with single-source batches, then 30K steps with mixed-source batches.
| Hyperparameter | Pre-Training | Fine-Tuning |
| Warmup ViT | 10% | 10% |
| Warmup Con. | 1% | 10% |
| Warmup LLM | 10% | 10% |
| LR ViT | 6e-6 | 5e-6 |
| LR Con. | 2e-4 | 5e-6 |
| LR LLM | 2e-5 | 1e-5 |
| Cosine Decay | 0.1 | 0.1 |
| Eps. | 1e-6 | 1e-6 |
| Betas | 0.9, 0.95 | 0.9, 0.95 |
| Batch Size | 128 | 256 |
| Steps | 25K | 60K |
| Samples | 3.2M | 15.3M |
| Tokens | 10B | 37B |
| GPU Hours | 296 | 1,000 |
5 Evaluation
We compare jina-vlm against lightweight VLMs across six capability areas: general VQA, multimodal comprehension, multi-image reasoning, hallucination control, mathematical reasoning, text-only performance, and multilingual understanding. All evaluations use VLMEvalKit222https://github.com/open-compass/VLMEvalKit (vlmevalkit) with English prompts matching our training format (e.g., “Return only the letter of the best answer option” for multiple-choice, “Respond very briefly” for open-ended questions).
5.1 General VQA Tasks
Table 3 reports results on eight VQA benchmarks covering diagrams (AI2D (ai2d)), charts (ChartQA (chartqa), CharXiv (charxiv)), scene text (TextVQA (textvqa)), documents (DocVQA (docvqa), InfoVQA (infovqa)), OCR (OCRBench (ocrbench)), and diverse scenes (SEED-Bench-2-Plus (seedbench2plus)). jina-vlm achieves the highest average (72.3), with particularly strong performance on diagram interpretation and text extraction.
| Model Name | AI2D | ChartQA | TextVQA | DocVQA | InfoVQA | OCR | SEED-2 | CharXiv | Overall |
| (test avg) | (val) | (val) | (val) | Bench | Plus | (RQ / DQ) | |||
| jina-vlm | 32.3 / 63.5 | ||||||||
| Qwen2-VL-2B | * | * | 23.3 / 55.0* | ||||||
| Qwen3-VL-2B | * | * | * | 28.8 / 62.3 | |||||
| InternVL3-2B | * | * | 28.3 / 54.7 | ||||||
| InternVL3.5-2B | * | * | 31.6 / 65.0 |
Results for models other than jina-vlm are from their respective papers, internvl_3_5; internvl_3; qwen_2_vl, except those marked with * which were computed using VLMEvalKit. All scores represent accuracy (%) except OCRBench which uses a 0–1000 scale; for overall average computation, OCRBench scores are divided by 10 to align with the 0–100 scale of other benchmarks.
5.2 Document and Real-World Understanding
Table 4 shows results on multimodal comprehension (MME (mme), MMB v1.1 (mmbench), MMStar (mmstar)) and real-world understanding (RealWorldQA (realworldqa), MME-RealWorld (mmerealworld), R-Bench (rbench)). jina-vlm scores 67.4 on multimodal tasks and 61.9 on real-world tasks, achieving the best RealWorldQA result (68.2).
| Model | MME | MMB v1.1 | MMStar | Overall | RealWorld | MME-RW | R-Bench | Overall |
| (sum) | (EN) | (MM) | QA | (EN) | (dis) | (RW) | ||
| jina-vlm | ||||||||
| Qwen2-VL-2B | * | * | ||||||
| Qwen3-VL-2B | * | * | * | |||||
| InternVL3-2B | ||||||||
| InternVL3.5-2B |
Results for models other than jina-vlm are from their respective papers, internvl_3_5; internvl_3; qwen_2_vl, except those marked with * which are computed using VLMEvalKit. All scores represent accuracy (%) except MME which uses a 0–2800 scale; for overall average computation, MME scores are divided by 28 to align with the 0–100 scale of other benchmarks.
5.3 Multi-Image Reasoning and Hallucination
Table 5 reports multi-image reasoning (BLINK (blink), MuirBench (muirbench), MMT (mmtbench)) and hallucination benchmarks that measure the tendency to fabricate visual details (HallBench (hallusionbench), POPE (pope)). jina-vlm scores 47.3 on multi-image tasks, which is expected given limited multi-image training data, but achieves the best POPE score (90.3), indicating low hallucination rates.
| Model | BLINK | Muir | MMT | Overall | HallBench | POPE | Overall |
| (val) | Bench | (val) | (MI) | (avg) | (avg) | (Hall) | |
| jina-vlm | |||||||
| Qwen2-VL-2B | * | * | |||||
| Qwen3-VL-2B | * | * | |||||
| InternVL3-2B | |||||||
| InternVL3.5-2B |
Results for models other than jina-vlm are from their respective papers, (internvl_3_5; internvl_3; qwen_2_vl), except those marked with * which are computed using VLMEvalKit. All scores represent accuracy (%).
5.4 Mathematical Reasoning
Table 6 reports structured reasoning benchmarks: multidisciplinary comprehension (MMMU (mmmu)), visual mathematics (MathVista (mathvista), MathVision (mathvision), MathVerse (mathverse), WeMath (wemath)), and logical reasoning (LogicVista (logicvista)). jina-vlm performs comparably to InternVL3-2B and outperforms Qwen2-VL-2B.
| Model | MMMU | MathVista | MathVision | MathVerse | WeMath | LogicVista | Overall |
| (Vision Only) | |||||||
| jina-vlm | 59.5 | 19.2 | 23.9 | 17.1 | 33.3 | ||
| Qwen2-VL-2B | 43.0 | 12.4 | 17.3* | 10.9* | 27.3* | ||
| Qwen3-VL-2B | 61.3 | 31.6 | 22.7* | 28.0* | 35.4* | ||
| InternVL3-2B | 57.0 | 21.7 | 25.3 | 22.4 | 36.9 | ||
| InternVL3.5-2B | 71.8 / 61.5† | 42.8 / 26.5† | 53.4 / 35.3† | 48.5 / 19.1† | 47.7 / 41.4† |
Results for models other than jina-vlm are from their respective papers, (internvl_3_5; internvl_3; qwen_2_vl), except those marked with * which are computed using VLMEvalKit. indicates scores for InternVL3.5-2B without thinking mode, evaluated using VLMEvalKit. All scores represent accuracy (%).
5.5 Text-Only Performance
Table 7 compares jina-vlm against the backbone Qwen3-1.7B on text-only benchmarks: MMLU (mmlu), MMLU-Pro (mmlupro), GSM-8K (gsm8k), ARC-C (arc), and HellaSwag (hellaswag). Results show mixed preservation of text-only capabilities: jina-vlm matches or exceeds the backbone on commonsense reasoning (ARC-C, HellaSwag) and retains most performance on MMLU and GSM-8K. However, MMLU-Pro shows substantial degradation (46.4 30.3), likely because this benchmark emphasizes extended multi-step reasoning that conflicts with our instruction-tuning toward concise visual responses. This suggests a trade-off between optimizing for multimodal tasks and preserving complex text-only reasoning, which future work could address through more balanced data mixtures or curriculum scheduling.
| Model | MMLU | MMLU-Pro | GSM-8k | ARC-C | HellaSwag | Overall |
| jina-vlm | ||||||
| Qwen3-1.7B |
Results are collected using our evaluation code. All scores represent accuracy (%).
5.6 Multilingual Understanding
Table 8 reports multilingual multimodal benchmarks: MMMB (mmmb), Multilingual MMBench (mmmb), and MTVQA (mtvqa). jina-vlm achieves state-of-the-art multilingual performance among 2B-scale VLMs, with the highest averages on MMMB (78.8) and Multilingual MMBench (74.3).
| Benchmark | jina-vlm | Qwen2-VL-2B | Qwen3-VL-2B | InternVL3-2B | InternVL3.5-2B | |
| MMMB | ar | * | ||||
| cn | * | |||||
| en | * | |||||
| pt | * | |||||
| ru | * | |||||
| tr | * | |||||
| avg | * | |||||
| Multi. MMBench | ar | * | ||||
| cn | * | |||||
| en | * | |||||
| pt | * | |||||
| ru | * | |||||
| tr | * | |||||
| avg | * | |||||
| MTVQA | * | |||||
| Overall | ||||||
Results for baseline models are derived from their original publications, (internvl_3_5; internvl_3; qwen_2_vl), except those marked with * which are computed using VLMEvalKit. All scores represent accuracy (%).
6 Conclusion
We presented jina-vlm, a 2.4B vision-language model that achieves state-of-the-art multilingual visual question answering among open 2B-scale VLMs. Our results demonstrate that small VLMs can attain strong cross-lingual visual understanding through careful architectural and training choices: attention-based token pooling reduces visual tokens by 4 while preserving spatial information, and incorporating text-only data during multimodal training mitigates the catastrophic forgetting typically observed in vision-adapted language models. On standard English VQA benchmarks, jina-vlm achieves leading results, demonstrating that multilingual capabilities need not come at the cost of general performance.
The current approach has limitations. Multi-tile processing introduces computational overhead that scales with image resolution, and tiling can fragment global spatial context, potentially impairing performance on tasks requiring holistic scene understanding such as object counting or precise spatial reasoning across tile boundaries. While the global thumbnail partially mitigates this, native-resolution approaches (navit) may be better suited for such tasks. We have not emphasized safety-critical training or alignment, and multi-image reasoning remains weak due to limited training data in this regime. Future work could explore more efficient resolution handling, targeted improvements for counting and spatial tasks, and investigate whether our multilingual training recipe transfers to larger model scales.
Appendix A Appendix
A.1 Pseudocode for Creating Overlapping Tiles
Base input size ()
Patch size ();
Maximum number of tiles ( by default, configurable)
Overlap margins in patches ()
Tiling = (number of rows, number of columns)